# Towards Lifelong Dialogue Agents via Relation-aware Memory Construction and Timeline-augmented Response Generation

### v2

@@ 6 Further Analyses and Discussions
A concern of Theanine is the API cost. Regardless, we argue that it is competitive when both performance and cost are taken into account. Figure 9 plots response quality (Mauve score) against the API cost. [^11] Calculated Based on session 5, which involves most memories for management. We use Mauve for its stronger correlation with humans over diverse metrics (Pillutla et al., 2021). We find Theanine and all ablations not only outperform all baselines but also lie on the Pareto frontier, indicating an efficient cost-performance trade-off. This suggests Theanine’s value when performance is prioritized over API costs.
@@
Another inquiry left might be whether the growing span of memory will eventually hinder memory retrieval in Theanine if there ever are hundreds of sessions. Although this can be a serious issue for conventional memory retrieval, we presume that it will be mitigated to some extent in Theanine because: (i) We retrieve relevant memories as a whole in the form of timelines. This serves as a safety net in scenarios where an important memory is missed out in top-$k$ retrieval – it might still be retrieved along with the timeline; (ii) We perform timeline refinements to further adjust retrieved timelines based on current dialogue context such that it provided tailored information for the response generator. This acts as a second insurance against sub-optimal retrieval.
@@ 7 Related Work
Since MSC, there are several studies on long-term conversations: Bae et al. (2022) train a classifier to update old memories in phone call scenarios. As we enter the era of LLMs, Wang et al. (2023) and Li et al. (2024) leverage LLMs to write and update memories for RG. Apart from LLMs’ power, human behaviors also foster methods in this field. For example, Zhong et al. (2024) apply humans’ forgetting curve to make memories that have been discussed stay in the memory pool longer. Recently, Maharana et al. (2024) also adopt the concept of timelines. However, in their work, a timeline is a fixed, pre-defined series of events (potentially unrelated), which simply serve as a user profile for synthesizing dialogue data. By contrast, in our work, a timeline is built with relevant events, which are dynamically linked and retrieved as the conversation goes on, serving as useful cues for response generation.
@@ 8 Conclusions
This paper presents a first-ever timeline-based framework for memory-augmented dialogue agents in long-term conversations. Comprehensive evaluations show that [Uncaptioned image]Theanine facilitates more beneficial augmentation, leading to responses that are closer to ground-truths and more aligned with speakers’ past interactions. Its effectiveness is further confirmed in TeaFarm, a QA-based pipeline we propose to address the limitation of G-Eval and human efforts in assessing memory-augmented RG. We expect our novel approach to serve as a new foundation towards lifelong dialogue agents.
9 Limitations
@@
10 Ethical Statements
LLMs might generate harmful, biased, offensive, sexual content. Authors avoid such content from appearing in this paper. We guarantee fair compensation for human evaluators from Amazon Mechanical Turk. We ensure an effective pay rate higher than 20\$ per hour based on the estimated time required to complete the tasks.
11 Acknowledgments
This work was partly supported by an IITP grant funded by the Korean Government (MSIT) (No. RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)). Jinyoung Yeo is the corresponding author (jinyeo@yonsei.ac.kr).
@@ Appendix A Appendix Contents
  • Appendix B.1: Cause-effect Commonsense Relations.
@@
  • Appendix F: Session-specific Evaluation Results.
@@
  • Appendix I: Further Analyses and Discussions.
@@ Appendix C TeaFarm Evaluation
The overview of TeaFarm is illustrated in Figure 10.
@@ Appendix G Empirical Examples
We provide several empirical examples of [Uncaptioned image]Theanine. Examples of relation-aware memory linking are in Figure 1516, and 17. Examples of utilizing refined timeline for response generation are in Figure 18.
@@ Appendix H Prompts
Appendix I Further Analyses and Discussions
@@
At the end of each session, we use ChatGPT (gpt-3.5-turbo-0125) to summarize the conversation into memory sentences. We conduct examinations on such conversation-memory summarization using 100 randomly sampled sessions from MSC and CC to make sure the quality of raw memories is acceptable. The result is in Table 4.

### v3

@@ 5.2 TeaFarm Results
To provide insight regarding conversation scenarios that are challenging for dialogue agents, we present case studies of how Theanine fail in TeaFarm in Appendix G.
@@ 6 Further Analyses and Discussions
A concern of Theanine is the API cost. Regardless, we argue that it is competitive when both performance and cost are taken into account. Figure 9 plots response quality (Mauve score) against the API cost. [^11] Calculated based on session 5, which involves most memories for management. We use Mauve for its stronger correlation with humans (Pillutla et al., 2021). We find Theanine and all ablations not only outperform all baselines but also lie on the Pareto frontier, indicating an efficient cost-performance trade-off. This suggests Theanine’s value when performance is prioritized over API costs. Actual API costs and results based on B-4, R-L, and Bert scores are in Appendix I.
@@
Time efficiency.
Time efficiency can be an important consideration when deploying Theanine to real-world scenarios having richer events. Figure 10 shows time-performance comparisons regarding both “memory construction” and “retrieval + RG” also using the Pareto frontier. Similarly, Theanine and many of its ablations demonstrate an efficient time-performance trade-off.
Figure 10: Time-performance comparisons.
Additional comparison: Memory Retrieval with a dynamically-changing k.
Due to Theanine’s graph-based procedures, the response generator may access different amounts of memories during RG depending on given contexts (i.e., queries used by the retriever) and when the conversation takes place (i.e., an earlier or a later session), whereas conventional methods (Xu et al., 2022a; Bae et al., 2022) often have a fixed number $k$ of memories retrieved for RG. Therefore, to further quantify the effect of our proposed timeline-based management and augmentation, we compare Theanine to Memory Retrieval with a dynamic $k$, where k dynamically changes based on the number of collected memories in Theanine for each specific test data. In other words, if Theanine uses timelines to collect $k$ memories during RG for a test instance $\mathcal{D}_{i}$, baselines will also be retrieving $k$ memories for generating a response for $\mathcal{D}_{i}$.
Methods / Metrics Bleu-4 Rouge-L Mauve Bert
Memory Retrieval (dynamic k) 3.06 17.97 33.33 87.32
+ Memory Update 2.68 17.19 28.49 87.11
[Uncaptioned image]Theanine (Ours) 4.22 19.22 45.53 87.70
Table 5: Additional comparison, where $k$ in Memory Retrieval is dynamically modified for each test instance.
In Table 5, we can observe that when the number of memories is matched, ours outperforms both baselines despite the same amount of memories being provided. We assume this is because: (i) our graph-based retrieval helps us collect more beneficial memories than conventional retrieval; (ii) addressing the relation between events and shaping them based on dialogue contexts can facilitate richer contextual cues for RG.
@@
Another inquiry is whether the growing span of memory will eventually hinder retrieval in Theanine if there ever are hundreds of sessions. Although this may be a serious issue for conventional methods, we presume that it will be partially mitigated in Theanine, as: (i) We retrieve relevant memories as a whole in the form of timelines. This serves as a safety net in scenarios where an important memory is missed out in top-$k$ retrieval–it may be collected via the linked structure; (ii) We refine retrieved timelines based on current dialogue such that they provide tailored information for RG. This acts as a second insurance against sub-optimal retrieval.
@@ 7 Related Work
Since MSC, there have been several studies on long-term conversations: Bae et al. (2022) train a classifier to update old memories in phone call scenarios. As we enter the era of LLMs, Li et al. (2024) leverages LLMs to write and update memories for RG. Apart from LLMs’ power, human behaviors also foster methods in this field. For example, Zhong et al. (2024) apply humans’ forgetting curve to make memories that have been discussed exist longer. Recently, Park et al. (2023) and Maharana et al. (2024) also adopt the concept of timelines. However, Park et al. (2023) focus on tagging the timestamp (e.g., “22:00”) of events and does not explicitly model the connection between them, and, in Maharana et al. (2024), a timeline is a fixed, pre-defined series of events (potentially unrelated) which simply serve as a user profile for synthesizing dialogue data. By contrast, in our work, a timeline is built with relevant events, which are dynamically linked based on their causal relations and retrieved as the conversation goes on, benefitting our goal of consistent memory tracking and integration.
@@ 8 Conclusions
This paper presents the first-ever timeline-based memory management and augmentation framework, [Uncaptioned image]Theanine, for autonomous agents in long-term conversations. Applying Theanine, we develop a dialogue agent that efficiently addresses the constant, lifelong tracking of memories and their integration for response generation throughout its service life. Comprehensive evaluations show that [Uncaptioned image]Theanine can facilitate more beneficial memory augmentation, leading to responses that are closer to ground truths and more aligned with speakers’ past interactions. Theanine’s effectiveness is further confirmed in TeaFarm, a counterfactual-driven pipeline we design to address the limitation of G-Eval and human efforts in assessing memory augmentation. We expect our novel approaches to serve as a new foundation for future efforts towards lifelong dialogue agents.
Limitations
@@
Ethical Statements
LLMs might generate harmful, biased, offensive, and sexual content. Authors avoid such content from appearing in this paper. We guarantee fair compensation for human evaluators from Amazon Mechanical Turk. We ensure an effective pay rate higher than 20\$ per hour based on the estimated time required to complete the tasks.
Acknowledgments
This work was mainly supported by STEAM R&D Project, NRF, Korea (RS-2024-00454458) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project), and was partially supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2024-00333484; RS-2024-00414981). Jinyoung Yeo is the corresponding author (jinyeo@yonsei.ac.kr).
@@ Appendix A Appendix Contents
  • Appendix B.1: Cause-effect Commonsense Relations Adopted.
@@
  • Appendix F: Session-specific Results of Automatic Evaluation.
@@
  • Appendix I: Further Analyses.
@@ Appendix C TeaFarm Evaluation
The overall pipeline of TeaFarm is illustrated in Figure 11.
@@ Appendix G Empirical Examples
Outputs from Theanine.
We provide several empirical examples of [Uncaptioned image]Theanine. Examples of relation-aware memory linking are in Figure 1617, and 18. Examples of utilizing refined timeline for response generation are in Figure 19.
How Theanine fails in TeaFarm.
We present failure cases where Theanine fails to pass the TeaFarm test in Figure 20 and Figure 21. In Figure 20, although the conversation has shifted to “librarian”, the similarity-based retriever retrieves unhelpful memories due to the huge portion of “kid” in the context. While a helpful memory (i.e., “A is a retired libraria”) is eventually caught by our designed timeline structure, the LLM still hallucinate. We assume it is due to the noises introduced by those highly-ranked, yet irrelevant memories, and it highlights the need for addressing helpfulness ranking among retrieved memories in lifelong dialogue systems. Figure 21 shows a failure case, where Theanine successfully retrieves the correct memories but generates an improper response. We hypothesize that this is because relation-aware linking and context-aware timeline refinement may sometimes make the length of input tokens too long such that the agent cannot properly utilize key information provided. We believe this can be resolved to an extent via dedicated prompt (i.e., the prompt for RG) engineering. We leave this to future work.
@@ Appendix H Prompts
Appendix I Further Analyses
@@
At the end of each session, we use ChatGPT (gpt-3.5-turbo-0125) to summarize the conversation into memory sentences. We conduct examinations on such summarization using 100 randomly sampled sessions from MSC and CC to make sure the quality of raw memories is acceptable. The result is in Table 6.
@@
Cost-efficiency trade-off assessed using other metrics.
In Section 6, we have presented methods having an efficient cost-performance trade-off (i.e., are Pareto-efficient) by plotting the Mauve score against API cost (Figure 9). We present methods that are Pareto-efficient when considering the other three metrics used in our study, i.e., B-4, R-L, and Bert Score, in Table 7.
Agents B-4 R-L Bert Score
All Dialogue History
All Memories
+ Update
Memory Retrieval $\checkmark$ $\checkmark$
+ Update
Rsum-LLM
MemoChat
COMEDY
Theanine (ours) $\checkmark$ $\checkmark$
w/o Relation-aware Linking
w/o Refinement $\checkmark$
Shuffled Timeline $\checkmark$ $\checkmark$ $\checkmark$
Table 7: Methods considered Pareto-efficient when judged based on B-4, R-L, and Bert Score reported in Table 1. $\checkmark$ = Pareto-efficient methods.
API costs.
The actual API costs of all settings (ours and baselines) are in Table 8.
Agents Cost Ratio (ours $=$ 1) Cost (per episode; \$)
All Dialogue History 0.50 0.0067
All Memories & $\mathcal{D}$ 0.27 0.0036
+ Update 5.71 0.0771
Memory Retrieval 0.17 0.0023
+ Update 5.63 0.0760
Rsum-LLM 0.42 0.0057
MemoChat 0.52 0.0076
COMEDY 0.61 0.0082
Theanine (ours) 1.00 0.0135
w/o Relation-aware Linking 0.50 0.0067
w/o Refinement 0.71 0.0096
Shuffled Timeline 0.20 0.0027
Table 8: API costs for Theanine and baselines.
@@
Figure 20: Theanine fails to pass TeaFarm (Example 1) - Due to sudden topic change.
Figure 21: Theanine fails to pass TeaFarm (Example 2) - Due to sub-optimal timeline utilization during RG.